OpenStudy (anonymous):

Let R^3 have the Euclidean inner product, and suppose that u =(1,1,1) v =(6,7,-15.Find the value of k for which ||ku+v||=13

OpenStudy (anonymous):

any idea, friends, I got negative value for k^2. stuck!!!

OpenStudy (anonymous):

Why don't you brute force it?

OpenStudy (anonymous):

what you mean?

OpenStudy (anonymous):

\[ ||ku+v||=13\\ ||ku+v||^2=13^2\\ <ku+v,ku+v>=13^2 \]

OpenStudy (anonymous):

so, it is not ||ku+v||^2 = ||ku||^2 +||v||^2= k^2||u||^2 +\\v||^2 =13^2?

OpenStudy (anonymous):

I think they want you to use inner products here.

OpenStudy (anonymous):

ohoh, see, can you see my weak logic at breaking down the ||ku||^2

OpenStudy (anonymous):

Also: \[ ||a+b|| \neq ||a||+||b|| \]

OpenStudy (anonymous):

yes, you are right, but square is another topic.

OpenStudy (anonymous):

Consider if \(a=-b\)

OpenStudy (anonymous):

yes, you are right, my bad. if they are orthogonal, then sqr is applied

OpenStudy (anonymous):

ku = (k,k,k) , right? break down follow that way, right?

OpenStudy (anonymous):

thanks a lot, friend, I think I can get the answer from your way. got it, k != negative value anymore.